# IterPref: Focal Preference Learning for Code Generation via Iterative Debugging

### v2

@@ 7 Ablation Study
Ref. 83.5 78.7 79.4 69.0 77.7
SFT 87.8 83.5 82.3 69.6 80.8
Hybrid-DPO 83.5 79.3 81.5 66.1 77.9
Hybrid-RPO 84.8 79.3 81.5 67.7 78.8
DiffAug-DPO 83.5 76.8 80.2 65.1 76.4
DiffAug-RPO 86.0 79.9 81.5 65.9 78.8
IterPref-DPO 89.6 85.4 83.9 69.8 82.2
IterPref-RPO 89.6 86.0 82.5 70.4 82.1
@@
Ref. 61.6 53.0 76.9 62.9 63.6
SFT 87.2 82.9 83.1 68.3 80.4
Hybrid-RPO 82.9 79.3 81.7 67.5 77.9
DiffAug-RPO 86.0 81.7 82.8 67.5 79.5
IterPref-RPO 89.6 84.8 83.3 69.5 81.9
Table 2: Pass@1 score (%) when performing supervised fine-tuning and possible ablations on the IterPref.
Symmetric Masking Strategy. If training with the symmetric masking strategy, where Code LLMs learn from both $D^{+}$ and $D^{-}$ without access to the full positive sample, the model fails to retain its basic code generation capabilities and cannot benchmark properly. The primary objective of Code LLMs is to generate complete and correct code. Although learning symmetrically from both $D^{+}$ and $D^{-}$ may seem appealing, the priority is ensuring that Code LLMs learn from complete, correct code rather than fragmented pieces. Without the full correct code contexts, the positive sample cannot align with the instruction, introducing incomplete and misleading signals into the learning process.
8 Conclusion
We present IterPref, a novel preference alignment framework that emulates human iterative debugging to capture critical errors in code for precise optimization. IterPref leverages preference pairs to identify error-prone regions and applies an improved DPO algorithm, guiding Code LLMs to focus on aligning these pivotal segments. This targeted approach enhances the model’s ability to detect and correct mistakes more efficiently. To facilitate the generation of high-quality preference pairs, we contribute the CodeFlow training set, where each sample undergoes iterative refinement until it passes unit tests, with the modification history providing a natural record of error corrections. Extensive experiments show that IterPref-equipped Code LLMs achieve significant performance improvements in code generation and excel in tackling basic and complex coding tasks like BigCodeBench.
Limitations
IterPref is inspired by the debugging pattern of developers, serving as a novel framework for fine-grained preference learning in Code LLMs. Instead of using pass rate alone, IterPref derives preference pairs from an iterative debugging process. By contrasting critical tokens between a corrected version and its preceding iteration, IterPref helps the model learn fine-grained alignment for key errors. The limitations for the IterPref framework are twofold: first, IterPref relies on generated test cases, and when the quality of these test cases is not ensured, the performance of IterPref may also be affected. Second, this study focuses on a dataset of 59k samples without further expansion, which may limit generalizability, but offers opportunities for future exploration with larger data set.

### v3

@@ B.2 Comparison with Supervised Fine-Tuning
Ref. 83.5 78.7 79.4 69.0 39.6 43.6 15.3 58.4
SFT 87.8 83.5 82.3 69.6 35.9 45.6 17.0 60.2
Target-DPO 89.6 85.4 83.9 69.8 39.9 48.7 20.2 62.5
Target-RPO 89.6 86.0 82.5 70.4 38.3 48.4 19.9 62.2
@@
Ref. 61.6 53.0 76.9 62.9 40.2 45.8 24.1 52.1
SFT 87.2 82.9 83.1 68.3 39.1 51.6 30.0 63.2
DiffAug-RPO 86.0 81.7 82.8 67.5 40.7 51.4 30.5 62.9
Target-RPO 89.6 84.8 83.3 69.5 43.1 53.3 33.3 65.3
Table 9: SFT and our results on CodeQwen1.5-7B-Chat and Qwen2.5-Coder-7B. Detailed results on LiveCodeBench and BigCodeBench are presented in Table 10 and Table 11.
Table 10 provides a more granular breakdown of performance on LiveCodeBench-v5 by difficulty, where our method consistently outperforms the SFT versions of both CodeQwen-7B-Chat and Qwen-Coder-7B, especially on Easy and Medium problems.
EASY MEDIUM HARD Avg
CodeQwen-7B-Chat-SFT 41.87 9.14 0.75 17.06
CodeQwen-7B-Chat-Ours 48.73 11.89 0.75 20.16
Qwen-Coder-7B-SFT 67.14 21.03 2.61 30.01
Qwen-Coder-7B-Ours 69.31 27.13 3.73 33.33
Table 10: Detailed Results on LiveCodeBench-v5, comparing SFT with Our Target-DPO.
Similarly, Table 11 shows detailed results on BigCodeBench categories, again illustrating the advantage of our approach over SFT across different task types and difficulties. For example, Qwen-Coder-7B with our method achieves an average of 36.8 compared to SFT’s 31.8.
Complete-Full Instruct-Full Complete-Hard Instruct-Hard Average
CodeQwen-7B-Chat-SFT 45.6 35.9 18.3 15.5 28.8
CodeQwen-7B-Chat-Ours 48.4 38.3 20.3 18.2 31.3
Qwen-Coder-7B-SFT 51.6 39.1 21.6 14.9 31.8
Qwen-Coder-7B-Ours 53.3 43.1 29.7 20.9 36.8
Table 11: Detailed Results on BigCodeBench, comparing SFT with Our Target-DPO.
B.3 Comparison with Other Code Generation Methods
We further benchmark our method against other notable code generation techniques. Table 12 benchmarks our method against techniques like CodeDPO, Code-Optimise, and PLUM on HumanEval and MBPP. Notably, on the DeepSeekCoder-6.7B base, our method achieves significantly higher scores (e.g., 66.50 on HumanEval vs. 59.75 for CodeDPO and 56.70 for PLUM).
HumanEval HumanEval+ MBPP MBPP+
StarCoder2-7B 35.40 29.90 54.40 45.60
CodeDPO 48.17 34.15 58.40 49.37
Code-Optimise 32.32 28.05 58.90 47.89
PLUM 46.30 39.60 60.40 49.10
Our Target-DPO 48.20 43.90 63.50 50.60
DeepSeekCoder-1.3B 31.53 28.65 57.40 48.67
CodeDPO 42.07 38.04 61.37 53.43
Code-Optimise 34.15 30.49 59.15 49.87
Our Target-DPO 47.00 43.30 61.37 54.20
DeepSeekCoder-6.7B 47.60 39.60 70.20 56.60
CodeDPO 59.75 51.83 72.18 60.01
Code-Optimise 47.56 37.20 72.18 57.64
PLUM 56.70 48.80 72.90 58.90
Our Target-DPO 66.50 60.40 76.50 61.40
Table 12: Performance comparison with baselines.
Appendix C In-depth Analyses and Ablation Studies
In this section, we conduct several in-depth analyses and ablation studies to better understand the characteristics and behavior of our proposed method. This includes investigating the impact of model size (scaling laws), sensitivity to key hyperparameters ($\beta$ and $\alpha$), the diversity of our constructed preference data, common error patterns in the generated code, and the efficiency of our preference annotation process.
C.1 Scaling Law on Model Size
The impact of model size on performance when applying our method is detailed in Table 13. The results show a clear trend: as model size increases from 1.5B to 32B parameters, the average performance improves from 54.4 to 74.7, demonstrating the scalability of our approach.
Model HumanEval MBPP BCB-Inst BCB-Comp Average
Base Plus Base Plus
Qwen2.5-Coder-1.5B 67.7 62.8 66.7 55.3 33.4 40.5 54.4
Qwen2.5-Coder-7B 89.6 84.8 83.3 69.5 43.1 53.3 70.6
Qwen2.5-Coder-32B 92.7 86.6 89.4 74.6 45.7 58.9 74.7
Table 13: Ablations on model size.
C.2 Ablation Studies on Hyperparameters $\beta$ and $\alpha$
In Direct Preference Optimization (DPO), the hyperparameter $\beta$ controls the strength of the preference signal, essentially determining how strictly the model should adhere to the learned preferences relative to the reference model. The hyperparameter $\alpha$, when part of the DPO framework or a combined loss, often serves as a weighting factor for an additional objective or regularization term. We performed ablation studies on key hyperparameters $\beta$ and $\alpha$. Table 14 presents the results for $\beta$ when $\alpha$ is set to 0, suggesting that a smaller $\beta$ (e.g., 0.1) yields the best average performance (69.7).
$\beta$ HumanEval MBPP BCB-Inst BCB-Comp Average
Base Plus Base Plus
0.1 89.0 83.6 83.1 69.0 41.0 52.7 69.7
0.3 86.6 80.5 82.5 68.5 40.3 50.3 69.1
0.5 85.4 80.5 83.6 66.7 40.5 48.2 67.5
Table 14: Ablations on $\beta$ with $\alpha$ set 0.
The corresponding ablation for $\alpha$, with $\beta$ fixed at 0.1, is shown in Table 15. These results indicate that $\alpha=1.0$ provides the highest average score (70.5), while $\alpha=\infty$ (equivalent to SFT) performs relatively worse.
$\alpha$ HumanEval MBPP BCB-Inst BCB-Comp Average
Base Plus Base Plus
1.0 89.6 84.8 83.1 69.3 43.1 53.3 70.5
3.0 87.2 81.1 82.8 69.0 40.3 52.7 67.9
5.0 84.1 80.5 83.1 70.1 40.3 54.0 68.7
$\infty$ (SFT) 87.2 82.9 83.1 68.3 39.1 51.6 68.7
Table 15: Ablations on $\alpha$ with $\beta$ set 0.1.
C.3 Ablation on Context Usage
Beyond CodeQwen1.5-7B-Chat, we also provide ablation results on the use of context from the rejected sample using Qwen2.5-Coder-7B, as shown in Table 16.
HumanEval MBPP BCB-Inst BCB-Comp LCB-v5 Average
Base Plus Base Plus
Qwen2.5-Coder-7B 61.6 53.0 76.9 62.9 40.2 45.8 24.1 52.1
SFT 87.2 82.9 83.1 68.3 39.1 51.6 30.0 63.2
Hybrid-RPO 82.9 79.3 81.7 67.5 41.2 50.5 29.8 61.8
DiffAug-RPO 86.0 81.7 82.8 67.5 40.7 51.4 30.5 62.9
Target-RPO 89.6 84.8 83.3 69.5 43.1 53.3 33.3 65.3
Table 16: Ablation results on the Target-DPO using Qwen2.5-Coder-7B.
C.4 Analysis of Preference Data Diversity
To understand the characteristics of our CodeFlow preference dataset, Table 17 provides a distributional analysis of various features (e.g., Workflow, Functionality, Data Processing) across 1k samples, comparing it with other common datasets like Alpaca and OSS-Instruct. Our CodeFlow dataset (both preferred and dis-preferred samples) generally exhibits a higher count and thus potentially greater diversity across most features, particularly in Data Processing, File Operation, and Advanced Techniques.
Datasets Workflow Functionality
Computation
Operation
User
Interaction
Data
Processing
File
Operation
Alpaca 994 393 282 82 221 11
CodeFeedback 2079 535 689 143 895 39
Evol-Alpaca 2163 591 783 134 1401 55
OSS-Instruct 2254 669 413 192 903 102
CodeFlow (Preferred) 2689 805 967 410 2418 290
CodeFlow (Dis-Preferred) 2490 772 964 406 2327 287
Logging Algorithm
Data
Structures
Implementation
Logic
Advanced
Techniques
Average
Alpaca 1 232 72 67 10 215.00
CodeFeedback 10 427 100 49 63 457.18
Evol-Alpaca 15 414 130 74 94 532.18
OSS-Instruct 62 150 140 82 26 453.91
CodeFlow (Preferred) 133 790 367 152 178 836.27
CodeFlow (Dis-Preferred) 129 785 361 149 193 805.73
Table 17: Distribution of total features across 1k samples.
C.5 Error Analysis of Generated Code
Target-DPO Generates Fewer Errors.
In this section, we present a statistical analysis of common failure case types to pinpoint frequent pitfalls in code generation. By contrasting critical tokens between a corrected version and its preceding iteration explicitly, a Code LLM equipped with Target-DPO makes fewer errors. Table 18 presents the frequency of common failure types (e.g., AttributeError, KeyError) on the BigCodeBench Complete-Full set. Our Target-RPO method shows a notable reduction in the sum of these errors (308 occurrences) compared to RPO (369) and Code-Optimise-RPO (396) on Qwen2.5-Coder-7B. This suggests that while RPO includes SFT, it still requires targeted learning of critical errors in the dispreferred samples to effectively reduce mistakes.
Error Type RPO Code-Optimise-RPO Target-RPO (Ours)
AttributeError: ‘X’ has no attribute ‘Y’ 145 149 127
KeyError: ‘X’ 139 128 100
NameError: name ‘X’ is not defined 41 58 46
FileNotFoundError: No such file or directory 44 61 35
Sum 369 396 308
Table 18: The frequency of the most common failure types on the BigCodeBench Complete-Full set.
C.6 Efficiency Analysis of Preference Annotation
Additionally, we compare the costs of generating and annotating preference pairs to guide more efficient preference alignment and reduce errors.
Target-DPO Provides an Efficient Pathway for Preference Annotation.
We compare the cost of synthesizing one preference pair between Target-DPO and the sampling techniques adopted by Code-Optimise, primarily considering external LLM calls and execution times. Given an instruction, Code-Optimise synthesizes $m$ code snippet candidates (where $m$ is often set to 100), using $n$ test cases from the raw dataset, leading to $m\times n$ executions on the CPU. In contrast, for a single instruction, Target-DPO requires up to 7 LLM calls and executions for successful pair generation in most cases. Considering the failure ratio (when code can’t pass the generated test cases within the budget), an estimated 10.4 calls are needed for a given instruction on average across the dataset. This is far fewer than the $m$ (e.g., 100) calls for sampling candidates plus subsequent executions often employed by sampling-heavy methods. Though massive sampling can yield diverse candidates, it is not efficient as most code snippets are discarded. Target-DPO shows that starting with a single code snippet, even if it fails initially, it still holds high potential to form a valuable preference pair for alignment training through iterative refinement.
Method HE HE+ MBPP MBPP+ BCB-Comp BCB-Inst LCB-v5 Avg.
Ref. 61.6 53.0 76.9 62.9 45.8 40.2 24.1 52.1
Prefix-Suffix 86.6 79.9 81.5 68.0 47.8 39.2 30.1 61.9
LCS 89.0 83.6 83.1 69.0 52.7 41.0 32.6 64.4
Table 19: Comparison of Prefix-Suffix and LCS methods for code difference extraction across benchmarks.
Method HE HE+ MBPP MBPP+ BCB-Comp BCB-Inst LCB-v5 Avg.
Ref. 61.6 53.0 76.9 62.9 45.8 40.2 24.1 52.1
First 87.8 81.7 81.0 67.5 47.2 40.4 29.7 62.2
Last 88.4 82.9 82.0 68.0 51.7 41.2 31.5 63.7
Random (Ours) 89.0 83.6 83.1 69.0 52.7 41.0 32.6 64.4
Table 20: Comparison of different negative sample selection strategies across benchmarks.
Method HE HE+ MBPP MBPP+ BCB-Comp BCB-Inst LCB-v5 Avg.
Ref. 61.6 53.0 76.9 62.9 45.8 40.2 24.1 52.1
DS-v3-SFT 82.3 76.2 83.3 69.8 46.9 40.7 26.4 60.8
DS-v3-Ours 84.8 79.3 84.7 72.0 47.5 40.7 27.7 62.4
Table 21: Results of distillation from DeepSeek-V3 to Qwen2.5-Coder-7B.
C.7 Alternatives to LCS for Code Differencing
To explore alternatives beyond LCS-based code differencing, we considered a simple baseline approach where we extract the common prefix and suffix, treating the middle part as the difference. The results are shown in Table 19:
Consider these two versions of a Python function where only two lines are truly different:
Listing 1: Chosen Code
def hello():
print("Hello") # <-- difference
x = 1 # <-- unchanged
y = 2 # <-- unchanged
return x + y # <-- difference
z = hello()
Listing 2: Rejected Code
def hello():
print("Hi") # <-- difference
x = 1 # <-- unchanged
y = 2 # <-- unchanged
return x * y # <-- difference
z = hello()
Prefix-Suffix-based method incorrectly marks both x = 1 and y = 2 as differences, while the LCS-based method correctly identifies only the two truly changed lines (print("Hi") and return x * y, providing a more precise grounding of the differences, enabling more targeted training.
C.8 Different Negative Samples
We chose the random negative selection method to better simulate the variety of error types in real-world scenarios, where some versions are severely flawed and others only slightly. This random slection approach enhances the diversity of negative samples and helps improve generalization.
To validate this choice, we conduct additional experiments using only the first or last incorrect versions as negatives. The results in Table 20 show that our method consistently achieves superior performance on six benchmarks.
C.9 Different Teacher Model
We adopt DeepSeek-V3 (671B MoE LLM with 37B active parameters) as the teacher model and synthesize a total of 28k pairs with resource constraints. The results using Qwen2.5-Coder-7B as the student model are shown in Table 21. Target-DPO demonstrates generalizable improvement when using different teachers for distillation.